OpenStudy (anonymous):

find dy/dx y=sin(x)+(1/2)cot(x) don't give me the answer. i want to work through with it

sam (.sam.):

Pretty straight forward

sam (.sam.):

What have you did?

OpenStudy (anonymous):

[sin (x)]((1/2)cot(x))+(sin (x))[1/2)cot(x)]

OpenStudy (anonymous):

(cos(x))(1/2)cot(x))+(sin (x))((1/2)csc^2(x))

OpenStudy (anonymous):

what do i do after that?

OpenStudy (anonymous):

That is wrong

OpenStudy (anonymous):

What is the dy/dx of sin, and the dy/dx of cot

OpenStudy (anonymous):

sin(x) is cos(x) and cot(x) is csc^2(x) right?

OpenStudy (anonymous):

First u divided 2 part dy/dx or sin = cos(x) right, then next part u take dx/dx or 1/2cot(x)

OpenStudy (anonymous):

-csc^2

OpenStudy (anonymous):

So the answer is be y= cos (x) - csc^2 (x) I guess

OpenStudy (anonymous):

u mean Y= cos (x) - (1/2)csc^2 (x)